a sum of $2800 is invested, partly at 5% and partly at 4%. if the total interest is $1280 per annum, find the amount of each investment.
Answer by ikleyn(52781)

.
a sum of $2800 is invested, partly at 5% and partly at 4%. if the total interest is $1280 per annum, find the amount of each investment.
~~~~~~~~~~~~~~


As the problem is posed,  IT  IS  INCORRECT  and has no solution.

Notice that even  5%  of the total amount of  2800  dollars,    0.05*2800 = 5*28 = 140  dollars  is   L E S S   than  $1280.


Thus the problem is a  FAKE.